OpenStudy (mayaal):

verify the identity: cot x/1+csc x=csc x-1/cot x

OpenStudy (neer2890):

\[\frac{ \cot x }{ 1+cosec x }= cosec x- \frac{ 1 }{ \cos x }??\]

OpenStudy (mayaal):

no,the right side is =(csc x-1)/cot x

OpenStudy (mayaal):

can someone help me,plz?

OpenStudy (neer2890):

\[\frac{ cotx }{ 1+cosecx }*\frac{ (1-cosec x) }{ (1-cosec x) }\] \[=\frac{ cotx *(1-cosec x) }{ (1-cosec ^{2}x) }\]

OpenStudy (neer2890):

also, we know that \[1+ \cot ^{2}x=cosec ^{2}x\] \[1- cosec ^{2}x= - \cot ^{2}x\] putting this value in above equation. \[\frac{ \cot x*(1-cosecx) }{ - \cot ^{2}x }= \frac{ -(1-cosecx )}{ \cot x }=\frac{ cosecx -1 }{ \cot x }\]

OpenStudy (mayaal):

thnku sooooooo much @neer2890 and @RaphaelFilgueiras :)

OpenStudy (neer2890):

you're welcome........:)

OpenStudy (anonymous):

you are welcome